# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v2

@@ 1. Introduction
We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [91].
@@
We prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (c) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect. $8$ of [11] that has not emerged in other parts.
@@
More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (c). This allows us to apply our abstract uniqueness Theorem 8.19 in [11] for braided tensor structures in a pre-tensor category.
@@ 2. The main result
    The representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ admits a natural structure of ribbon braided tensor category obtained by transporting Huang-Lepowsky structure to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$. In this way, Zhu’s linear equivalence
    $Z:{\rm Rep}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$
@@
    Let $A(V_{{\mathfrak{g}}_{k}})$ be endowed with the twisted unitary compatible coboundary weak quasi-Hopf structure induced by the weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ via Wenzl path and the construction of the Drinfeld twist, following Theorem 31.1 in [11] (a), and let ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
    Then ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way with the structure derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and the equivalence of Theorem 31.1 in [11],
    ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q)).$
    Moreover the transport in (a) preserves the tensor products, the ribbon for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples $(V_{\lambda},V,V)$, $(V,V_{\lambda},V)$,$(V,V,V_{\lambda})$ with $V_{\lambda}$ an arbitrary irreducible of the fusion category.
@@
    If ${\mathfrak{g}}$ is of one of the Lie types $A$, $B$, $C$, $D$, $G_{2}$ then the transport preserves the braided tensor structure for all objects. Thus $Z$ is a ribbon braided tensor equivalence. Specifically, the composition
@@
The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects. 29, 30 in [11], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [91] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix. The former more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
We would like to notice that the passages described in part (c) of Theorem 2.1 are explicit. The first and the last equivalences are due respectively, to the work by Zhu combined with that of Huang and Lepowsky. The last is the functor due to Wenzl enriched with our weak tensor structure, preserving Drinfeld coboundary structure derived from the ribbon and $R$-matrix, see the above cited sections in [11].
@@
The weak Hopf algebras constructed in [10] have been extended to all Lie types in Sections 29, 30 of [11] with uniform methods for all the Lie types.
@@
As said, on the the side of quantum groups, our approach has been influenced by results in quantum group theory by Woronowicz but also by the dual viewpoint by Majid for quasi-Hopf algebras, extended to weak quasi-Hopf algebras by Haring-Oldenburg. We start reviewing these Tannakian reconstruction theorems from the viewpoint of Majid and Haring-Oldenburg in Sect. 5 in [11], where the reader may find indication of the original references. Moreover Kazhdan and Wenzl theory is also important to us, and the reader may consult Sect. 23 in [11] and references therein. It is by no means an exaggeration to assert that our paper [11] would have not existed without the paper by Neshveyev, Tuset on Kazhdan-Lusztig theorem for compact quantum groups [71]. This paper was strongly inspiring for our strategy, although methods are different. For similar reasons, we were also guided by the presentation of Drinfeld-Kohno theorem in the book by Bakalov and Kirillov [1]. These last two references first inspired our work in the search of a unified treatment following a Tannakian approach in the idea of Majid, in our case. An important aspect in our case as far as Majid Tannakian duality for quasi-Hopf algebras, is the search of natural tensor structures for the involved embedding functors into the category of Hilbert spaces (Wenzl functor and Zhu functor in the two settings).
@@
The mathematical contribution by Tsuchiya and Kanie [88], and the subsequent works by A. Wassermann [90] and Toledano-Laredo [87] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [91] that led to our approach. The reader may find reference to more of the several original papers in mathematics and physics in Sections 31–36 in [11].
@@
We briefly explain here the plan of our proof of the main theorem for parts (a), (b), and (c). In the following sections we describe the main ideas to prove these parts in greater detail.
@@
For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [23], [25], [26], and in our case for different categories, defined in the statement of the main result. This is the main part of the result. The first main step is the construction of a class of finite dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ for all Lie types and positive integer levels, extending the former examples of [10] in the type $A$ case. They have unitary coboundary structures associated to the fusion category of quantum groups at roots of unity, associated to a specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [91]. Complete definitions of unitary coboundary structures on weak quasi-Hopf algebras are given in [11], and are motivated by the work of several authors, among them Drinfeld, Wenzl, Mack and Schomerus. See Sections 3, 5, 9, 10, 11, 26, 28 in [11].
In Sect. 8 of [11] we prove two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object. These are Theorem 8.13 and its extension Theorem 8.19. We explain here briefly how Theorem 8.19 applies to imply part (c) of Theorem 2.1.
The proof of parts (b) and (c) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 3–6, and 8–16 in [11]. This abstract theory splits in two parts. Theory on which part (b) of Theorem 2.1 is based is developed in Sections 3–6, and 9–16. The applicative part of the proof of part (b) is developed in Sections 19–36. The abstract aspects that lead to the proof of part (c) of Theorem 2.1 are the main subject of Section 8.
For part (b), we develop an abstract analogue of Drinfeld-Kohno theorem, Theorem 28.4, and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels. We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 30, by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [91] for the special tensor products described in (b). The weak tensor structure is constituted by fusion intertwiners with respect to the action of $U_{q}({\mathfrak{g}})$.
@@
The first way is traditionally referred to as Tannaka-Krein duality. One needs a (weak, quasi-) tensor structure $(F,G)$ on ${\mathcal{F}}$, and then apply Theorem 6.6 to obtain a (weak, quasi)-bialgebra structure on $A={\rm Nat}_{0}({\mathcal{F}})$ (with extra structure) with the desired equivalence property. The (weak quasi-) tensor structure
@@
In Sect. 30 we shall construct a weak tensor structure on Wenzl functor, and derive a weak Hopf finite dimensional $C^{*}$-algebra $A_{W}({\mathfrak{g}},q)$ associated to fusion categories of quantum groups at roots of unity for all Lie types, with extra structure, that is the braiding, unitary structures given by a Drinfeld coboundary matrix, also rigid and ribbon, extending the work of [10] for ${\mathfrak{s}l}_{N}$ in turn extending and expanding the work by Mack and Schomerus for ${\mathfrak{s}l}_{2}$ [64].
@@
We discuss this transfer in Sect. 15 without the need of a weak quasi-tensor structure on ${\mathcal{F}}$, in abstract terms. One only needs an inverse linear equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$. In this situation we may define a braided tensor category structure in ${\rm Rep}(A)$ from the braided tensor category structure in ${\mathcal{C}}$ with the property that ${\rm Rep}(A)$ is equivalent to ${\mathcal{C}}$ as a braided tensor category. The tensor product, associativity morphisms, braiding morphisms of ${\rm Rep}(A)$ are explicitly defined via the equivalence ${\mathcal{E}}$ and its inverse ${\mathcal{S}}$.
@@
Section 15 abstracts Zhu linear equivalence and its inverse, by considering a pair of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$ associated to ${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$ and a linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ and describes how to make ${\rm Rep}(A)$ into a braided tensor category.
@@
A central method of proof for both parts (b) and (c) is the use of the fundamental generating representation $V$ and its properties, following primarily the work of Wenzl. The paper [91] plays a main role especially for the proof of (b). For part (c) we shall later describe results and related works in the literature on which our proof is based.
Part (b) of Theorem 2.1 gives coincidence of the two braided tensor structures for all Lie types (two pairs of associativity morphisms together with their braiding morphisms satisfying the two hexagonal diagrams) on a specific collection of their variables, in the two settings of quantum groups and vertex operator algebras. These collections do not generate additively all the variables of the braiding morphisms or associativity morphisms, because they do not contain all the irreducible objects.
@@
The twist takes one unitary structure of a special tensor product arising in the quantum group setting from [91] to the classical unitary structure corresponding to the compact real form of ${\mathfrak{g}}$ (and this in turn remarkably leads to the construction of a unitary tensor structure on the module category of the vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ in a uniform way). This implies that the twist leaves the corresponding structural properties invariant on these special variables. To see this, we work on the side of the vertex operator algebras and we compare with the work by Frenkel and Zhu [35]. An abstract concept describing these common structural properties of the associativity morphisms is introduced in Sect. 8 with the name of a ${\mathcal{V}}$-pre-associator of CFT type. The name is motivated by our understanding of the structure of Huang-Lepowsky associativity morphisms.
@@
Direct verification of coincidence of the two braided tensor structures on all the irreducible objects is more complicated due to multiplicity of fusion and non-semisimplicity of $U_{q}({\mathfrak{g}})$. To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 8.13 and its generalization Theorem 8.19. This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the comparison between special cases of the two associativity morphisms and their corresponding braided symmetries. Thus the uniqueness theorem 8.19 is formulated with suitable assumptions, where again in the application the abstract notion of ${\mathcal{V}}$-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem. But this brief description of Theorems 8.13 and Theorem 8.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), is still not enough to describe our method of proof to complete the proof of part (c) and with it the proof of Theorem 2.1.
@@
We need more that we describe in the following sections. The first next step is a somewhat complicated cohomological reduction work done in Theorems 8.13 and 8.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
@@
The cohomological reduction is useful to reduce the problem to the last step, which uncovers a central duality property between the fundamental representation of the quantum group and corresponding representations of the braid group in the intertwining spaces of the fundamental representation, in the various Lie types. The first example of this is Schur-Weyl duality in the type $A$ case, and works in the literature that we shall recall greatly extend this first classical example to to other Lie types and to the quantum cases. We shall come back to this, but let us anticipate similarities and differences with the original Drinfeld-Kohno theorem, useful for the proof of (b), that will also useful to clarify how we were led to uncover the key role of duality of the fundamental representation $V$, for the proof of the last step (c).
@@
8. Main tools to prove (b): Analogies and differences between Drinfeld-Kohno Theorem and Theorem 28.4 in [11]
@@
We next outline general ideas that make our strategy possible, useful to prove part (b) of Theorem 2.1.
@@
Secondly, in our approach the notion of coboundary matrix first introduced by Drinfeld in his work on Drinfeld-Kohno theorem for quasi-Hopf algebras in a formal setting, plays a major two-fold role. In our case we adopt the idea of Drinfeld coboundary that we describe for complex weak quasi-Hopf algebras in an analytic setting. This general theory is considered in Sections 26, 27, 28.
@@
Similarities with the original Drinfeld-Kohno theorem also concern the identification method between the two associativity morphisms, in our case the twisted structure of that arising from quantum groups and that coming from Huang Lepowsky tensor structure both on the Zhu algebra. In Proposition 3.16 of [25], Drinfeld identifies the most difficult parts of the corresponding strictures for Drinfeld category using his uniqueness Theorem 3.15 based on cohomology of the same paper. In our case, we similarly develop a uniqueness theorem but for the associativity morphisms which we verify of our weak Hopf algebras $A_{W}({\mathfrak{g},q})$ and with very different methods, based on the generating representation $V$ for which we develop cohomological reduction Theorem 8.19 in [11] and extension to quantum case and different Lie types of Schur-Weyl duality. We discuss this more in the next two sections.
9. Differences between the original Drinfeld-Kohno theorem and our analogue Theorem 28.4 in the treatment of the associativity morphisms, to prove (c).
@@
Differently from the original Drinfeld-Kohno theorem, we do not study the problem of equivalence associativity morphisms in our version version of Drinfeld-Kohno theorem. We do this in Section 8 in [11] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
@@
10. Main tools to prove (c): cohomological reduction of a braided tensor structure of a category with a generating object $V$ to braid group duality of $V$

### v4

@@ 1. Introduction
We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [99]. On the other hand, the remaining Lie types may be either derived from Finkelberg equivalence, making now the ribbon braided fusion category $\tilde{\mathcal{O}}_{k}$ unitary, or treated with our partial results in the setting of Huang-Lepowsky structure for representation categories of vertex operator algebras. Direct constructions of unitary structures in this setting have been obtained by Gui as already mentioned [43], [44], [45], [46].
@@
As already mentioned, we prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (d) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect. 9 of [13] that has not emerged in other parts.
@@
More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (d). This allows us to apply our abstract uniqueness Theorem 9.19 in [13] for braided tensor structures in a pre-tensor category.
@@ 2. The main result
    The linear representation category ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ admits a natural structure of ribbon braided tensor category ${\rm Rep}_{\rm HL}(A(V_{{\mathfrak{g}}_{k}}))$ obtained as a transport from Huang-Lepowsky structure on ${\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})$ to ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$. In this way, Zhu’s linear equivalence
    $Z:{\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}_{\rm HL}(A(V_{{\mathfrak{g}}_{k}}))$
@@
    The Zhu algebra $A(V_{{\mathfrak{g}}_{k}})$ becomes a twisted unitary compatible coboundary weak quasi-Hopf algebra with structure induced by the weak Hopf algebra $A_{W}({\mathfrak{g}},q)$ via Wenzl path and the construction of the Drinfeld twist. Let ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ be endowed with the corresponding ribbon braided tensor structure.
    At the level of representation categories, ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ becomes a unitary coboundary modular tensor category in this way with the structure derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and the equivalence
    ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q)),$
    obtained in Theorem 3.2 in [13]. Moreover ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ and ${\rm Rep}_{\rm HL}(A(V_{{\mathfrak{g}}_{k}}))$ have the same tensor product bifunctor.
@@
    The linear equivalence
    $Z:{\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\to{\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$
    endowed with the same structure maps as in (a) but w.r.t. the new ribbon braided tensor structure of ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ described in (b), satisfies the ribbon braided tensor equivalence equations for the ribbon structure for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples
    $(V_{\lambda},V,V),\quad(V,V_{\lambda},V),\quad(V,V,V_{\lambda})$
    with $V_{\lambda}$ an arbitrary irreducible object of the fusion category.
  • (d)
    If ${\mathfrak{g}}$ is of one of the Lie types $A$, $B$, $C$, $D$, $G_{2}$ then Zhu’s equivalence in (c) is a ribbon braided tensor equivalence. Specifically, the composition
@@
The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects. 31, 32 in [13], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [99] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix. The more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
For part (b), we refer the reader to Theorem 3.2 (a) in [13] and further indications therein where the statement is explained.
We would like to notice that the passages described in (2.1) are explicit. The first and the last equivalences are due respectively, to the work by Zhu combined with that of Huang and Lepowsky. The last is the functor due to Wenzl enriched with our weak tensor structure, preserving Drinfeld coboundary structure derived from the ribbon and $R$-matrix, see the above cited sections in [13].
@@
The weak Hopf algebras constructed in [12] have been extended to all Lie types in Sections 31, 32 of [13] with uniform methods for all the Lie types.
@@
As said, on the the side of quantum groups, our approach has been influenced by results in quantum group theory by Woronowicz but also by the dual viewpoint by Majid for quasi-Hopf algebras, extended to weak quasi-Hopf algebras by Haring-Oldenburg. We start reviewing these Tannakian reconstruction theorems from the viewpoint of Majid and Haring-Oldenburg in Sect. 6, 7, in [13], where the reader may find indication of the original references. Moreover Kazhdan and Wenzl theory is also important to us, and the reader may consult Sect. 25 in [13] and references therein. It is by no means an exaggeration to assert that our paper [13] would have not existed without the paper by Neshveyev, Tuset on Kazhdan-Lusztig theorem for compact quantum groups [78]. This paper was strongly inspiring for our strategy, although methods are different. For similar reasons, we were also guided by the presentation of Drinfeld-Kohno theorem in the book by Bakalov and Kirillov [1]. These last two references first inspired our work in the search of a unified treatment following a Tannakian approach in the idea of Majid, in our case. An important aspect in our case as far as Majid Tannakian duality for quasi-Hopf algebras, is the search of natural tensor structures for the involved embedding functors into the category of Hilbert spaces (Wenzl functor and Zhu functor in the two settings).
@@
The mathematical contribution by Tsuchiya and Kanie [95], and the subsequent works by A. Wassermann [98] and Toledano-Laredo [94] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [99] that led to our approach. The reader may find reference to more of the several original papers in mathematics and physics in Sections 34–39 in [13].
@@
We briefly explain here the plan of our proof of the main theorem for parts (a), (b), (c), and (d). In the following sections we describe the main ideas to prove these parts in greater detail.
@@
For part (b), we follow a suitable Drinfeld twist method originally used by Drinfeld for Drinfeld-Kohno theorem [25], [27], [28], and in our case for different categories, defined in the statement of the main result. This is the main part of the result. The first main step is the construction of a class of finite dimensional weak Hopf $C^{*}$-algebras $A_{W}({\mathfrak{g}},q)$ for all Lie types and positive integer levels, extending the former examples of [12] in the type $A$ case. They have unitary coboundary structures associated to the fusion category of quantum groups at roots of unity, associated to a specific linear fiber functor $W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Vec}$, that emerges from the work of Wenzl [99]. Complete definitions of unitary coboundary structures on weak quasi-Hopf algebras are given in [13], and are motivated by the work of several authors, among them Drinfeld, Wenzl, Mack and Schomerus.
In Sect. 9 of [13] we prove two main abstract uniqueness results for associativity morphisms and braided symmetries in a semisimple pre-tensor category with a generating object. These are Theorem 9.13 and its extension Theorem 9.19. We explain here briefly how Theorem 9.19 applies to imply part (d) of Theorem 2.1.
The proof of parts (b), (c), and (d) is based on the general theory of semisimple weak quasi-Hopf algebras, weak Hopf algebras, and the theory of Drinfeld twist treated in Sections 5–8, and 10–18 in [13]. This abstract theory splits in two parts. Theory on which part (b), (c) of Theorem 2.1 is based is developed in Sections 5–8, and 10–17. The applicative part of the proof of part (b), (c) is developed in Sections 20–39. The abstract aspects that lead to the proof of part (d) of Theorem 2.1 are the main subject of Section 9.
For parts (b), (c) we develop an abstract analogue of Drinfeld-Kohno theorem 30.4 in [13], and apply it to the setting of fusion categories of quantum groups at suitable roots of unity and of module categories of affine vertex operator algebras at positive integer levels. We construct weak Hopf algebras $A_{W}({\mathfrak{g}},q)$ associated to the quantum group fusion category for all the Lie types in Section 32 in [13], by constructing a weak tensor structure on a linear functor from this category to the category of vector spaces, that emerges in [99] for the special tensor products described in (b), (c). The weak tensor structure is constituted by fusion intertwiners with respect to the action of $U_{q}({\mathfrak{g}})$.
@@
The first way is traditionally referred to as Tannaka-Krein duality. One needs a (weak, quasi-) tensor structure $(F,G)$ on ${\mathcal{F}}$, and then apply Theorem 8.6 in [13] to obtain a (weak, quasi)-bialgebra structure on $A={\rm Nat}_{0}({\mathcal{F}})$ (with extra structure) with the desired equivalence property. The (weak quasi-) tensor structure
@@
In Sect. 32 we shall construct a weak tensor structure on Wenzl functor, and derive a weak Hopf finite dimensional $C^{*}$-algebra $A_{W}({\mathfrak{g}},q)$ associated to fusion categories of quantum groups at roots of unity for all Lie types, with extra structure, that is the braiding, unitary structures given by a Drinfeld coboundary matrix, also rigid and ribbon, extending the work of [12] for ${\mathfrak{s}l}_{N}$ in turn extending and expanding the work by Mack and Schomerus for ${\mathfrak{s}l}_{2}$ [71].
@@
We discuss this transfer in Sect. 16 in [13] without the need of a weak quasi-tensor structure on ${\mathcal{F}}$, in abstract terms. One only needs an inverse linear equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$. In this situation we may define a braided tensor category structure in ${\rm Rep}(A)$ from the braided tensor category structure in ${\mathcal{C}}$ with the property that ${\rm Rep}(A)$ is equivalent to ${\mathcal{C}}$ as a braided tensor category. The tensor product, associativity morphisms, braiding morphisms of ${\rm Rep}(A)$ are explicitly defined via the equivalence ${\mathcal{E}}$ and its inverse ${\mathcal{S}}$.
@@
Section 16 in [13] abstracts Zhu linear equivalence and its inverse, by considering a pair of the Tannakian linear equivalence ${\mathcal{E}}:{\mathcal{C}}\to{\rm Rep}(A)$ associated to ${\mathcal{F}}:{\mathcal{C}}\to{\rm Vec}$ and a linear inverse equivalence ${\mathcal{S}}:{\rm Rep}(A)\to{\mathcal{C}}$ and describes how to make ${\rm Rep}(A)$ into a braided tensor category.
@@
A central method of proof for parts (b), (c), and (d) is the use of the fundamental generating representation $V$ and its properties, following primarily the work of Wenzl. The paper [99] plays a main role especially for the proof of (b), (c). For part (d) we shall later describe results and related works in the literature on which our proof is based.
Parts (b), (c) of Theorem 2.1 gives coincidence of the two braided tensor structures for all Lie types (two pairs of associativity morphisms together with their braiding morphisms satisfying the two hexagonal diagrams) on a specific collection of their variables, in the two settings of quantum groups and vertex operator algebras. These collections do not generate additively all the variables of the braiding morphisms or associativity morphisms, because they do not contain all the irreducible objects.
@@
The twist takes one unitary structure of a special tensor product arising in the quantum group setting from [99] to the classical unitary structure corresponding to the compact real form of ${\mathfrak{g}}$ (and this in turn remarkably leads to the construction of a unitary tensor structure on the module category of the vertex operator algebra $V_{{\mathfrak{g}}_{k}}$ in a uniform way). This implies that the twist leaves the corresponding structural properties invariant on these special variables. To see this, we work on the side of the vertex operator algebras and we compare with the work by Frenkel and Zhu [37]. An abstract concept describing these common structural properties of the associativity morphisms is introduced in Sect. 9 with the name of a ${\mathcal{V}}$-pre-associator of CFT type. The name is motivated by our understanding of the structure of Huang-Lepowsky associativity morphisms.
@@
Direct verification of coincidence of the two braided tensor structures on all the irreducible objects is more complicated due to multiplicity of fusion and non-semisimplicity of $U_{q}({\mathfrak{g}})$. To this aim rather than pursuing a closer analogy with the original Drinfeld-Kohno theorem for all objects, we develop two uniqueness theorems, Theorem 9.13 and its generalization Theorem 9.19. This generalization is made necessary by the actual conclusions reached in our work in the applicative sections on the comparison between special cases of the two associativity morphisms and their corresponding braided symmetries. Thus the uniqueness theorem 9.19 is formulated with suitable assumptions, where again in the application the abstract notion of ${\mathcal{V}}$-pre-associator of CFT-type for the fundamental representation $V$ is used in a central way both on the side of quantum groups and that of vertex operator algebras to verify coincidence of these special braid and associativity morphisms, after application of our Drinfeld-Kohno theorem. But this brief description of Theorems 9.13 and Theorem 9.19 put together with coincidence of the two braided tensor structures on special collections of objects proved in part (b), (c), is still not enough to describe our method of proof to complete the proof of part (d) and with it the proof of Theorem 2.1.
@@
We need more that we describe in the following sections. The first next step is a somewhat complicated cohomological reduction work done in Theorems 9.13 and 9.19, where we use cohomology of the pentagon equation and its interplay with the hexagonal equations defining the braided symmetry for a general semisimple braided tensor category endowed with a generating object.
@@
The cohomological reduction is useful to reduce the problem to the last step, which uncovers a central duality property between the fundamental representation of the quantum group and corresponding representations of the braid group in the intertwining spaces of the fundamental representation, in the various Lie types. The first example of this is Schur-Weyl duality in the type $A$ case, and works in the literature that we shall recall greatly extend this first classical example to to other Lie types and to the quantum cases. We shall come back to this, but let us anticipate similarities and differences with the original Drinfeld-Kohno theorem, useful for the proof of (b), (c), that will also useful to clarify how we were led to uncover the key role of duality of the fundamental representation $V$, for the proof of the last step (d).
@@
8. Main tools to prove (b), (c): Analogies and differences between Drinfeld-Kohno Theorem and Theorem 30.4 in [13]
@@
We next outline general ideas that make our strategy possible, useful to prove part (b), (c) of Theorem 2.1.
@@
Secondly, in our approach the notion of coboundary matrix first introduced by Drinfeld in his work on Drinfeld-Kohno theorem for quasi-Hopf algebras in a formal setting, plays a major two-fold role. In our case we adopt the idea of Drinfeld coboundary that we describe for complex weak quasi-Hopf algebras in an analytic setting. This general theory is considered in Sections 27, 28, 29, 30.
@@
Similarities with the original Drinfeld-Kohno theorem also concern the identification method between the two associativity morphisms, in our case the twisted structure of that arising from quantum groups and that coming from Huang Lepowsky tensor structure both on the Zhu algebra. In Proposition 3.16 of [27], Drinfeld identifies the most difficult parts of the corresponding structures for Drinfeld category using his uniqueness Theorem 3.15 based on cohomology of the same paper. In our case, we similarly develop a uniqueness theorem but for the associativity morphisms which we verify of our weak Hopf algebras $A_{W}({\mathfrak{g},q})$ and with very different methods, based on the generating representation $V$ for which we develop cohomological reduction Theorem 9.19 in [13] and extension to quantum case and different Lie types of Schur-Weyl duality. We discuss this more in the next two sections.
9. Differences between the original Drinfeld-Kohno theorem and our analogue Theorem 30.4 in the treatment of the associativity morphisms, to prove (d).
@@
Differently from the original Drinfeld-Kohno theorem, we do not study the problem of equivalence associativity morphisms in our version version of Drinfeld-Kohno theorem. We do this in Sect. 9 in [13] with the mentioned different ideas that appeal to duality properties of the fundamental representation $V$ with respect to the representation of the braid group in the tensor powers of $V$.
@@
10. Main tools to prove (d): cohomological reduction of a braided tensor structure of a category with a generating object $V$ to braid group duality of $V$